# CARTE: pretraining and transfer for tabular learning

### v1

@@ A.1 Pretrained model of CARTE
The model specification and training details were largely referenced from the work of Devlin et al. (2019). We set $12$ attention layers, each consisting of $12$ multi-head attentions, and the hidden dimension was fixed to the same size as the inputs ($300$). The resulting model contains over $9.3$ million parameters. To run the pretraining, we selected $128$ entities with one additional positive, resulting in the batch size of $256$. The total number of steps for training was $1,000,000$, which approximately covers $40$ epochs with respect to YAGO entities. We use the AdamW optimizer accompanied by the cosine scheduler with $lr_{min}=5\times 10^{-6}$, $lr_{max}=1\times 10^{-4}$ and a warmup over the first $10,000$ steps. The dropout rate was fixed to $0.1$ and the gelu activation function was used.
@@ A.2 Details on experiment settings for downstream tasks
To evaluate the performances of baselines on single tables, we focused on the few-shot learning setting with the number of train data for each table varying from $32$, $64$, $128$, $256$, $512$, and $1,024$; the rest of the remaining data were set as the test set. To find the optimal hyperparameters of the baselines, $5$-fold cross-validation over $100$ random search iteration were carried out on all the comparing methods except for CARTE and TabPFN. For CARTE, the same $5$-fold cross-validation, but only the grid-search over the learning rate was conducted. For Tabpfn, we ran with the default values, as suggested in the paper. For detailed information the hyperparameter spaces of each method, please refer to the Hyperparmeter tuning paragraph below. The performance was recorded on $10$ different train/test splits, with the performance measure set as the $R^{2}$ score for regression tasks and the Area Under Receiver Operating Curve (AUROC) for classification tasks.
@@
The hyperparameter tuning was done using grid search for CARTE, as we only tune the learning rate, and with random search for the baselines as these come with more than two hyperparameters to tune. The hyperparmeter spaces for the HGB, Resnet, and MLP baselines are based on that used in Grinsztajn et al. (2022); for CatBoost we follow that used in the CatBoost paper (Dorogush et al., 2018). For the baselines in joint learning across multiple tables, we employ an additional hyperparameter ‘fraction source’, which denote the fraction of source data used for training. The tables below depict the hyperparameter spaces for each.
Parameter Grid
Learning rate [$2.5$, $5$, $7.5$] $\times$ [$1e^{-4}$, $1e^{-3}$]
Table 1: CARTE hyperparameter space
Parameter Distribution
Max depth UniformInt [$2$, $6$]
@@
Table 2: CatBoost hyperparameter space, adapted from Dorogush et al. (2018)
Parameter Distribution
Learning rate LogUniform [$1e^{-2}$, $10$]
@@
Table 3: HistGradientBoosting hyperparameter space, adapted from Grinsztajn et al. (2022)
Parameter Distribution
Num layers UniformInt [$1$, $8$]
@@
Table 4: Resnet hyperparameter space, adapted from Grinsztajn et al. (2022)
Parameter Distribution
Num layers UniformInt [$1$, $4$]
@@
Table 5: MLP hyperparameter space, adapted from Grinsztajn et al. (2022)
Parameter Distribution
Source fraction Uniform [00, $1$]
Table 6: Additional hyperparameter for the baselines in joint learning
@@ B.2 Datasets used
    Beer Ratings  [^7] https://www.kaggle.com/datasets/ruthgn/beer-profile-and-ratings-data-set: The dataset contains tasting profiles and consumer reviews for 3197 unique beers from 934 different breweries. The task is to predict overall review ratings of different beers.
@@
    Cardekho  [^10] https://www.kaggle.com/datasets/sukritchatterjee/used-cars-dataset-cardekho. This dataset contains information on used cars, with their listing price in the websit Cardekho. The task is to predict the price.
@@
    Chocolate Bar Ratings  [^11] https://www.kaggle.com/datasets/rtatman/chocolate-bar-ratings.Dataset containing information and expert rating on cocoa batches. The task is to predict the rating.
@@
    Clear Corpus (Crossley et al., 2023) [^12] https://www.commonlit.org/blog/introducing-the-clear-corpus-an-open-dataset-to-advance-research-28ff8cfea84a: Generic information about the reading passage excerpts for elementary school students. The task is to predict the readability of the excerpts. The text feature is the name of the book, not the excerpt.
@@
    Company Employees [^14] https://www.kaggle.com/peopledatalabssf/free-7-million-company-dataset: Information on companies with over $1,000$ employees. The task is to predict the number of employees of each company.
@@
    Employee remuneration and expenses earning over 75000  [^15] https://opendata.vancouver.ca/explore/dataset/employee-remuneration-and-expenses-earning-over-75000/information/?disjunctive.department&disjunctive.title Remuneration and expenses for employees earning over \$75,000 per year. The task is to predict the remuneration of employees.
@@
    Employee Salaries  [^16] https://openml.org/d/42125: Information on salaries for employees of the Montgomery County, MD. The task is to predict the current annual salary range of the employees.
@@
    Fifa22 Players  [^17] https://www.kaggle.com/datasets/joebeachcapital/fifa-players: Information on soccer players and their ability scores in Fifa22 game. The task is to predict the player’s wage.
@@
    Filmtv Movies  [^18] https://www.kaggle.com/datasets/stefanoleone992/filmtv-movies-dataset/data: Information of movies and ratings scraped from an Italian movie review website Filmtv Movies. The task is to predict the public vote on movies.
@@
    Journal Score JCR: Scientific journals and their descriptive features from Journal Citation Reports. The task is to predict the impact factors of the journals.
@@
    Journal Score SJR: Scientific journals and their descriptive features from Scimago journal rank. The task is to predict the H-index of journals.
@@
    Japanese Anime:  [^19] https://www.kaggle.com/datasets/dbdmobile/myanimelist-dataset: List of Japanese animes and their relevant information. The task is to predict score for the animes.
@@
    K-Drama:  [^20] https://www.kaggle.com/datasets/noorrizki/top-korean-drama-list-1500: List of korean drama and their basic information from mydramalist website. The task is to predict the score of the korean dramas.
@@
    Michelin:  [^21] https://www.kaggle.com/datasets/ngshiheng/michelin-guide-restaurants-2021: List of restaurants along with additional details curated from the Michelin Restaurants guide. The task is to predict the award of the restaurants.
@@
    ML/DS Salaries:  [^22] https://ai-jobs.net/salaries/download/salaries.csv salary and basic information of workers in machine learning and data science industry. The task is to predict the salary of workers.
@@
    Movie Revenues:  [^23] https://www.kaggle.com/rounakbanik/the-movies-dataset: Metadata of movies released on or before July 2017. The task is to predict the range of the box-office revenues.
@@
    Museums:  [^24] https://www.kaggle.com/datasets/markusschmitz/museums: General information on the US museums. The task is to predict the revenues across the museums.
@@
    Mydramalist:  [^25] https://www.kaggle.com/datasets/rajchinagundi/mydramalist-complete-dataset General information on Asian drama scaped from mydramalist website. The task is to predict the ratings of Asian dramas.
@@
    NBA Draft:  [^26] https://www.kaggle.com/datasets/mattop/nba-draft-basketball-player-data-19892021 Information on all NBA Draft picks from 1989-2021. The task is to predict the ‘value over replacement’ of players.
@@
    Prescription Drugs:  [^27] https://data.ca.gov/uk/dataset/prescription-drugs-introduced-to-market The data contains new prescription drugs introduced to market in California with a Wholesale Acquisition Cost (WAC) that exceeding Medicare Part D. The task is to predict WAC at introduction.
@@
    Ramen ratings [^28] https://www.kaggle.com/datasets/ankanhore545/top-ramen-ratings-2022. The dataset contains ratings and characteristics of various ramens produced from multiple countries. The task is to predict the range of ratings of the ramens.
@@
    Roger Ebert: [^29] https://github.com/gabrielcs/movie-ratings-prediction. The dataset contains movies ratings by famous critic Rogert Ebert. The task is to predict the range of ratings.
@@
    Rotten Tomatoes:(Sanjib et al., 2023)  [^30] http://pages.cs.wisc.edu/~anhai/data/784_data/movies1/csv_files/rotten_tomatoes.csv. Contain information on movies that can be found in Rotten Tomatoes movie rating website. The task is to predict the rating values of the movies.
@@
    Spotify [^31] https://www.kaggle.com/datasets/maharshipandya/-spotify-tracks-dataset: Generic information on Spotify tracks with some associated audio features. The task is to predict the popularity of the albums.
@@
    US Accidents [^32] https://smoosavi.org/datasets/us_accidents: Information of accidents in US cities between 2016 and 2020. From this dataset, two tasks are conducted: (1) the range of accident counts for the US cities (2) the severity of the reported accidents.
@@
    US Presidential (Cvetkov-Iliev et al., 2022): Voting statistics in the 2020 US presidential election along with information on US counties. The task is to predict the range of voting numbers across US counties.
@@
    Used Cars 24:  [^33] https://www.kaggle.com/datasets/avikasliwal/used-cars-price-prediction. Information on used cars. The task is to predict the price.
@@
    Used Cars Benz Italy:  [^34] https://www.kaggle.com/datasets/bogdansorin/second-hand-mercedes-benz-registered-2000-2023-ita Dataset containing information on used cars sold in Italy. The task is to predict the price.
@@
    UsedCars.com: Dataset containing information on used cars usedcars.com. The task is to predict the price.
@@
    Used Cars Pakistan:  [^35] https://www.kaggle.com/datasets/mustafaimam/used-car-prices-in-pakistan-2021. Dataset containing information on used cars sold in Pakistan. The task is to predict the price.
@@
    Used Cars Saudi Arabia:  [^36] https://www.kaggle.com/datasets/turkibintalib/saudi-arabia-used-cars-dataset Dataset containing information on used cars sold in Saudi Arabia from the Syarah Website. The task is to predict the price of used cars.
@@
    Videogame Sales:  [^37] https://www.kaggle.com/datasets/gregorut/videogamesales This dataset contains a list of video games with sales greater than 100,000 copies (scrape of vgchartz.com). The task is to predict the global sales of the videogames.
@@
    Whisky: [^38] https://whiskyanalysis.com/index.php/database/ Basic and tasting information on whiskies form the whiskyanalysis.com. The task is to predict the range of meta critic of the whiskies.
@@
    Wikiliq:  [^39] https://www.kaggle.com/datasets/limtis/wikiliq-dataset Information on alcohol that can be found in the Wikiliq website. We conducted two tasks to predict the prices of (1) beer and (2) spirits.
@@
    Wina Poland:  [^40] https://www.kaggle.com/datasets/skamlo/wine-price-on-polish-market Information about wines on the polish market. The task is to predict the price.
@@
    Wine.com: [^41] https://www.kaggle.com/datasets/manyregression/updated-wine-enthusiast-review Information on wines scraped from the wine.com website. We conducted two tasks on prediction of (1) wine ratings and (2) wine prices.
@@
    WineEnthusiasts:  [^42] https://www.kaggle.com/datasets/manyregression/updated-wine-enthusiast-review Information about a wine and a taster from winemag.com. We conducted two tasks on prediction of (1) wine ratings and (2) wine prices.
@@
    WineVivino:  [^43] https://www.kaggle.com/datasets/joshuakalobbowles/vivino-wine-data. Information about wine bottles scrapped from Vivino’s website. We conducted two tasks on prediction of (1) wine ratings and (2) wine prices.
@@
    Yelp:  [^44] https://www.yelp.com/dataset The Yelp Open dataset for academic research. We extracted information on the restaurants from the original dataset. The task is to predict the range of stars for the restaurants. https://www.yelp.com/dataset
@@
    Zomato [^45] https://www.kaggle.com/datasets/anas123siddiqui/zomato-database?select=restaurant.csv. Information and reviews of restaurants found in the zomato websites. The task is to predict the range of ratings for each restaurants.

### v2

@@ A.1 Pretrained Model of CARTE
The model specification and training details were largely referenced from the work of (Devlin et al., 2019). We set $12$ attention layers, each consisting of $12$ multi-head attentions, and the hidden dimension was fixed to the same size as the inputs ($300$). The resulting model contains over $9.3$ million parameters. To run the pretraining, we selected $128$ entities with one additional positive, resulting in the batch size of $256$. The total number of steps for training was $1,000,000$, which approximately covers $40$ epochs with respect to YAGO entities. We use the AdamW optimizer accompanied by the cosine scheduler with $lr_{min}=5\times 10^{-6}$, $lr_{max}=1\times 10^{-4}$ and a warmup over the first $10,000$ steps. The dropout rate was fixed to $0.1$ and the gelu activation function was used.
@@ A.2 Details on Experiment Settings for Downstream Tasks
To evaluate the performances of baselines on single tables, we focused on the setting with limited train-size for each table varying from $32$, $64$, $128$, $256$, $512$, $1,024$, and $2,048$; the rest of the remaining data were set as the test set. To find the optimal hyperparameters of the baselines, $5$-fold cross-validation over $100$ random search iteration were carried out on all the comparing methods except for CARTE and TabPFN. For CARTE, the same $5$-fold cross-validation, but only the grid-search over the learning rate was conducted. For TabPFN, we ran with the default values, as suggested in the paper. For detailed information the hyperparameter spaces of each method, please refer to the Hyperparmeter tuning paragraph below. The performance was recorded on $10$ different train/test splits, with the performance measure set as the $R^{2}$ score for regression and the Area Under Receiver Operating Curve (AUROC) for classification tasks.
@@
The hyperparameter tuning was done using grid search for CARTE, as we only tune the learning rate, and with random search for the baselines as these come with more than two hyperparameters to tune. The hyperparmeter spaces for XGBoost, HistGradientBoosting, RandomForest, Resnet, and MLP baselines are based on that used in Grinsztajn et al. (2022); for CatBoost we follow that used in the CatBoost paper (Dorogush et al., 2018). For the baselines in joint learning across multiple tables, we employ an additional hyperparameter ‘fraction source’, which denote the fraction of source data used for training. Table 4 below summarizes the hyperparameter spaces for each of the estimators.
Table 4: Hyperparameter space for CARTE and baseline estimators.
Methods Parameters Grid
CARTE Learning rate [$2.5$, $5$, $7.5$] $\times$ [$1e^{-4}$, $1e^{-3}$]
CatBoost Max depth UniformInt [$2$, $10$]
@@
XGBoost Num estimators UniformInt [$50$, $1000$]
Max depth UniformInt [$2$, $10$]
Learning rate LogUniform [$1e^{-5}$, $1$]
Min child weight LogUniform [$1$, $100$]
Subsample Uniform [$0.5$, $1$]
Colsample by level Uniform [$0.5$, $1$]
Colsample by tree Uniform [$0.5$, $1$]
Gamma LogUniform [$1e^{-8}$, $7$]
Lambda LogUniform [$1$, $4$]
Alpha LogUniform [$1e^{-8}$, $100$]
HistGradientBoosting Learning rate LogUniform [$1e^{-2}$, $10$]
@@
RandomForest Num estimators UniformInt [$50$, $250$]
Max depth [None, $2$, $3$, $4$]
Max features [sqrt, sqrt, log2, None, $0.1$, $0.2$, $0.3$, $0.4$, $0.5$, $0.6$, $0.7$, $0.8$, $0.9$]
Min samples leaf LogUniform [$1.5$, $50.5$]
Bootstrap [True, False]
Min impurity decrease [00, $0.01$, $0.02$, $0.05$]
ResNet Num layers UniformInt [$1$, $8$]
@@
MLP Num layers UniformInt [$1$, $4$]
@@
Ridge Regression Solver [svd, cholesky, lsqr, sag]
Alpha LogUniform [$1e^{-5}$, $100$]
Logistic Regression Solver [newton-cg, lbfgs, liblinear]
Penalty [none, $l_{1}$, $l_{2}$, elasticnet]
C LogUniform [$1e^{-5}$, $100$]
Baselines in joint learning Source fraction Uniform [00, $1$]
A.3 Hardware Specifications
The pretrained model for CARTE was trained on GPUs. For rest of our experiments, they were run on 32 cores of CPU and the hardware was chosen based on availability.
GPUs:
NVIDIA V100 (32GB VRAM)
CPUs:
AMD EPYC 7742 64-Core Processor, AMD EPYC 7702 64-Core Processor (512GB RAM), Intel(R) Xeon(R) CPU E5-2660 v2, Intel(R) Xeon(R) Gold 6226R CPU (256GB RAM)
A.4 Implementation of CARTE
The implementation of CARTE will be available at https://github.com/soda-inria/carte.
@@ B.2 Datasets
    Beer Ratings [^7] https://www.kaggle.com/datasets/ruthgn/beer-profile-and-ratings-data-set The dataset contains tasting profiles and consumer reviews for 3197 unique beers from 934 different breweries. The task is to predict overall review ratings of different beers.
@@
    Cardekho [^10] https://www.kaggle.com/datasets/sukritchatterjee/used-cars-dataset-cardekho This dataset contains information on used cars, with their listing price in the websit Cardekho. The task is to predict the price.
@@
    Chocolate Bar Ratings [^11] https://www.kaggle.com/datasets/rtatman/chocolate-bar-ratingsDataset containing information and expert rating on cocoa batches. The task is to predict the rating.
@@
    Clear Corpus (Crossley et al., 2023) [^12] https://www.commonlit.org/blog/introducing-the-clear-corpus-an-open-dataset-to-advance-research-28ff8cfea84a Generic information about the reading passage excerpts for elementary school students. The task is to predict the readability of the excerpts. The text feature is the name of the book, not the excerpt.
@@
    Company Employees [^14] https://www.kaggle.com/peopledatalabssf/free-7-million-company-dataset Information on companies with over $1,000$ employees. The task is to predict the number of employees of each company.
@@
    Employee remuneration and expenses earning over 75000151515https://opendata.vancouver.ca/explore/dataset/employee-remuneration-and-expenses-earning-over-75000/information/?disjunctive.department&disjunctive.title Remuneration and expenses for employees earning over \$75,000 per year. The task is to predict the remuneration of employees.
@@
    Employee Salaries [^16] https://openml.org/d/42125 Information on salaries for employees of the Montgomery County, MD. The task is to predict the current annual salary range of the employees.
@@
    Fifa22 Players [^17] https://www.kaggle.com/datasets/joebeachcapital/fifa-players Information on soccer players and their ability scores in Fifa22 game. The task is to predict the player’s wage.
@@
    Filmtv Movies [^18] https://www.kaggle.com/datasets/stefanoleone992/filmtv-movies-dataset/data Information of movies and ratings scraped from an Italian movie review website Filmtv Movies. The task is to predict the public vote on movies.
@@
    Journal Score JCR Scientific journals and their descriptive features from Journal Citation Reports. The task is to predict the impact factors of the journals.
@@
    Journal Score SJR Scientific journals and their descriptive features from Scimago journal rank. The task is to predict the H-index of journals.
@@
    Japanese Anime [^19] https://www.kaggle.com/datasets/dbdmobile/myanimelist-dataset List of Japanese animes and their relevant information. The task is to predict score for the animes.
@@
    K-Drama [^20] https://www.kaggle.com/datasets/noorrizki/top-korean-drama-list-1500 List of korean drama and their basic information from mydramalist website. The task is to predict the score of the Korean dramas.
@@
    Michelin [^21] https://www.kaggle.com/datasets/ngshiheng/michelin-guide-restaurants-2021 List of restaurants along with additional details curated from the Michelin Restaurants guide. The task is to predict the award of the restaurants.
@@
    ML/DS Salaries [^22] https://ai-jobs.net/salaries/download/salaries.csv salary and basic information of workers in machine learning and data science industry. The task is to predict the salary of workers.
@@
    Movie Revenues [^23] https://www.kaggle.com/rounakbanik/the-movies-dataset Metadata of movies released on or before July 2017. The task is to predict the range of the box-office revenues.
@@
    Museums [^24] https://www.kaggle.com/datasets/markusschmitz/museums General information on the US museums. The task is to predict the revenues across the museums.
@@
    Mydramalist [^25] https://www.kaggle.com/datasets/rajchinagundi/mydramalist-complete-dataset General information on Asian drama scraped from mydramalist website. The task is to predict the ratings of Asian dramas.
@@
    NBA Draft [^26] https://www.kaggle.com/datasets/mattop/nba-draft-basketball-player-data-19892021 Information on all NBA Draft picks from 1989-2021. The task is to predict the ‘value over replacement’ of players.
@@
    Prescription Drugs [^27] https://data.ca.gov/uk/dataset/prescription-drugs-introduced-to-market The data contains new prescription drugs introduced to market in California with a Wholesale Acquisition Cost (WAC) that exceeding Medicare Part D. The task is to predict WAC at introduction.
@@
    Ramen ratings [^28] https://www.kaggle.com/datasets/ankanhore545/top-ramen-ratings-2022 The dataset contains ratings and characteristics of various ramens produced from multiple countries. The task is to predict the range of ratings of the ramens.
@@
    Roger Ebert [^29] https://github.com/gabrielcs/movie-ratings-prediction The dataset contains movies ratings by famous critic Rogert Ebert. The task is to predict the range of ratings.
@@
    Rotten Tomatoes (Sanjib et al., 2023) [^30] http://pages.cs.wisc.edu/~anhai/data/784_data/movies1/csv_files/rotten_tomatoes.csv Contain information on movies that can be found in Rotten Tomatoes movie rating website. The task is to predict the rating values of the movies.
@@
    Spotify [^31] https://www.kaggle.com/datasets/maharshipandya/-spotify-tracks-dataset Generic information on Spotify tracks with some associated audio features. The task is to predict the popularity of the albums.
@@
    US Accidents [^32] https://smoosavi.org/datasets/us_accidents Information of accidents in US cities between 2016 and 2020. From this dataset, two tasks are conducted: (1) the range of accident counts for the US cities (2) the severity of the reported accidents.
@@
    US Presidential (Cvetkov-Iliev et al., 2023) Voting statistics in the 2020 US presidential election along with information on US counties. The task is to predict the range of voting numbers across US counties.
@@
    Used Cars 24333333https://www.kaggle.com/datasets/avikasliwal/used-cars-price-prediction Information on used cars. The task is to predict the price.
@@
    Used Cars Benz Italy [^34] https://www.kaggle.com/datasets/bogdansorin/second-hand-mercedes-benz-registered-2000-2023-ita Dataset containing information on used cars sold in Italy. The task is to predict the price.
@@
    UsedCars.com Dataset containing information on used cars usedcars.com. The task is to predict the price.
@@
    Used Cars Pakistan [^35] https://www.kaggle.com/datasets/mustafaimam/used-car-prices-in-pakistan-2021 Dataset containing information on used cars sold in Pakistan. The task is to predict the price.
@@
    Used Cars Saudi Arabia [^36] https://www.kaggle.com/datasets/turkibintalib/saudi-arabia-used-cars-dataset Dataset containing information on used cars sold in Saudi Arabia from the Syarah Website. The task is to predict the price of used cars.
@@
    Videogame Sales [^37] https://www.kaggle.com/datasets/gregorut/videogamesales This dataset contains a list of video games with sales greater than 100,000 copies (scrape of vgchartz.com). The task is to predict the global sales of the videogames.
@@
    Whisky [^38] https://whiskyanalysis.com/index.php/database/ Basic and tasting information on whiskies form the whiskyanalysis.com. The task is to predict the range of meta critic of the whiskies.
@@
    Wikiliq [^39] https://www.kaggle.com/datasets/limtis/wikiliq-dataset Information on alcohol that can be found in the Wikiliq website. We conducted two tasks to predict the prices of (1) beer and (2) spirits.
@@
    Wina Poland [^40] https://www.kaggle.com/datasets/skamlo/wine-price-on-polish-market Information about wines on the polish market. The task is to predict the price.
@@
    Wine.com [^41] https://www.kaggle.com/datasets/manyregression/updated-wine-enthusiast-review Information on wines scraped from the wine.com website. We conducted two tasks on prediction of (1) wine ratings and (2) wine prices.
@@
    WineEnthusiasts [^42] https://www.kaggle.com/datasets/manyregression/updated-wine-enthusiast-review Information about a wine and a taster from winemag.com. We conducted two tasks on prediction of (1) wine ratings and (2) wine prices.
@@
    WineVivino [^43] https://www.kaggle.com/datasets/joshuakalobbowles/vivino-wine-data Information about wine bottles scrapped from Vivino’s website. We conducted two tasks on prediction of (1) wine ratings and (2) wine prices.
@@
    Yelp [^44] https://www.yelp.com/dataset The Yelp Open dataset for academic research. We extracted information on the restaurants from the original dataset. The task is to predict the range of stars for the restaurants. https://www.yelp.com/dataset
@@
    Zomato [^45] https://www.kaggle.com/datasets/anas123siddiqui/zomato-database?select=restaurant.csv Information and reviews of restaurants found in the zomato websites. The task is to predict the range of ratings for each restaurants.